OpenStudy (anonymous):

If the density of a block of wood that measures 2.0 cm on each side is 30.0 g/cm3 what is its mass?

OpenStudy (anonymous):

240g

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

\[2cm^3 = 8cm^3\] \[8cm^3 \times 30g/cm^3 = 240 grams\]

OpenStudy (anonymous):

i mean \[(2cm)^3\]

OpenStudy (anonymous):

Oh i see...

OpenStudy (anonymous):

volume = L*W*H = 2*2*2 mass = density*volume = 8*30 = 240

OpenStudy (anonymous):

Thank You :)